TFTOPL(1) TFTOPL(1) NAME tftopl-convertfontmetric(tfm)filestopropertylists SYNOPSIS tftopl[-verbose][-charcode-format=format]tfmfilename[plfilename] DESCRIPTION Thetftoplprogramtranslatesa(program-oriented)fontmetricfiletoa(human-oriented)property listfile.Thus,youcaneditthecontentsoftheTFMfiles,ifthefontdesignerhasnotdonehisor herjobproperly,orifyou’reencounteringstrangedifficulties,orifyou’rejustcurious.Italso servesasaTFM-filevalidatingprogram,i.e.,ifnoerrormessagesaregiven,theinputfileiscor-rect. Theplfilenamemustbecomplete.Noaddingofdefaultextensionsorpathsearchingisdone.If theplfilenameisnotgiven,standardoutputisused.Thetfmfilenamemustalsobecomplete; the‘.tfm’isnotadded.Butpathsearchingisdone:tftoplusestheenvironmentvariableTEX-FONTS.Ifthisisnotset,itusesthesystemdefault: TEXFONTS:/usr/TeX/lib/tex/fonts//:., Seetex(1)forthedetailsofthesearching. OPTIONS Withoutthe-verboseoption,tftoploperatessilently.Withit,abannerandprogressreportare printedonstdout. Theargumentformatto-charcode-formatspecifieshowcharactercodesareoutputinthePLfile. Bydefault,onlylettersanddigitsareoutputusingtheCintegercode(i.e.,inASCII);theothers areoutputinoctal.(Unlessthefont’scodingschemestartswithTeXmathsyorTeXmathex, inwhichcaseallcharactercodesareoutputinoctal.)Ifformatisascii,allcharactercodesthat correspondtographiccharacters,besidestheleftandrightparentheses,areoutputinASCII. Finally,ifformatisoctal,allcharactercodesareoutputinoctal. Seetex(1)forthedetailsofspecifyingoptions. SEEALSO pltotf(1) DonaldKnuth,TeXware,TheMETAFONTbook AUTHORS DonaldKnuthwrotetheprogram.ItwaspublishedaspartoftheTeXwaretechnicalreport,availablefromtheTeXUsersGroup.HowardTrickeyandPavelCurtisoriginallyportedittoUnix. 7Mar92 1